Person filling out Report: Anselm Wiercioch, XO
Summary Title: Longer EVA, beginning greenhab setup
Mission Status: Active. Full crew functional.
Sol Activity Summary: EVA’d to Dinosaur Quarry, set up first conveyor
stage in greenhab, planted first 12 plants.
Look Ahead Plan: EVA tomorrow to follow up on initial Quarry recon
Anomalies in work: back side of helmet cracked slightly on rock.
Repaired with superglue and zip ties. Not majorly concerning, but
helmet henceforth moved to reserve use only.
Weather: High 42F, Low 18F, wind avg 3.5mph, gust N/A, humidity
17-41%, grey cloudy skies
Crew Physical Status: Slightly stir crazy, but ok.
EVA: Anselm, Geoff, and Alison to Dinosaur Quarry
Reports to be filed:
– Sol Summary
– Commander’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
– Wifi still intermittent.